40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15261     Entry time: Sat Mar 7 15:18:30 2020
Author: gautam 
Type: Update 
Category: SUS 
Subject: EQ tripped some suspensions 

An earthquake around 330 UTC (=730pm yesterday eve) tripped ITMX, ITMY and ETMX watchdogs. ITMX got stuck. I released the stuck optic and re-enabled the local damping loops just now.

Attachment 1: EQ_6Mar.png  99 kB  | Hide | Hide all
EQ_6Mar.png
ELOG V3.1.3-